OpenStudy (anonymous):

Find each quotient express answer in simple form 7 3/5 ÷ 8 11/15 explain

OpenStudy (anonymous):

????

OpenStudy (displayerror):

How would you start the problem? What would you do first?

OpenStudy (anonymous):

To be honest dont not

OpenStudy (displayerror):

You're given two mixed fractions \(7 \dfrac{3}{5}\) and \(8 \dfrac{11}{15}\). I would turn those two into improper fractions. How would you go about converting them to improper fractions?

OpenStudy (anonymous):

This problem is on my book said find each quotient express answer in its simple form

OpenStudy (displayerror):

Yeah, you want to find the quotient (what you get as a result of dividing). In order to divide two mixed fractions (what you are given), you have to first convert to improper fractions. Have you learned about this in class yet? Try looking in your book. Here's a hint for the first one: \(7 \dfrac{3}{5}\) is the same as \(\dfrac{38}{5}\)

OpenStudy (anonymous):

Ahhh ok thank u